# Kernel Banzhaf: A Fast and Robust Estimator for Banzhaf Values

### v1

@@ Abstract
Banzhaf values offer a simple and interpretable alternative to the widely-used Shapley values. We introduce Kernel Banzhaf, a novel algorithm inspired by KernelSHAP, that leverages an elegant connection between Banzhaf values and linear regression. Through extensive experiments on feature attribution tasks, we demonstrate that Kernel Banzhaf substantially outperforms other algorithms for estimating Banzhaf values in both sample efficiency and robustness to noise. Furthermore, we prove theoretical guarantees on the algorithm’s performance, establishing Kernel Banzhaf as a valuable tool for interpretable machine learning.
@@ 1 Introduction
The increasing complexity of AI models has intensified the challenges associated with model interpretability. Modern machine learning models, such as deep neural networks and complex ensemble methods, often operate as “opaque boxes.” This opacity makes it difficult for users to understand and trust model predictions, especially in decision-making scenarios like healthcare, finance, and legal applications, which require rigorous justifications. Thus, there is a pressing need for reliable explainability tools to bridge the gap between complex model behaviors and human understanding.
Among the various methods employed within explainable AI, game-theoretic approaches have gained prominence for quantifying the contribution of features in predictive modeling and enhancing model interpretability. While primarily associated with feature attribution (Lundberg & Lee, 2017; Karczmarz et al., 2022), these methods also contribute to broader machine learning tasks such as feature selection (Covert et al., 2020) and data valuation (Ghorbani & Zou, 2019; Wang & Jia, 2023). Such applications extend the utility of explainable AI, fostering greater trust in AI systems by providing insights beyond traditional explanations.
@@
Shapley values, rooted in cooperative game theory, provide a principled way to attribute the contribution of $n$ individual players to the overall outcome of a game (Shapley, 1953). In the context of feature attribution, each “player” is a feature and the “game” is defined by a set function that maps a subset of features to the prediction score of an AI model on that subset. For a given feature, the Shapley value quantifies the average marginal contribution of the feature on the model’s prediction, computed as the weighted average over all possible combinations of features included in the model (see e.g., Equation 1) (Lundberg & Lee, 2017).
@@
An alternative to Shapley values are Banzhaf values, which also compute each individual’s contribution to the system’s overall outcome (Banzhaf, 1965). While Shapley values are more widely used, Banzhaf values are often considered more intuitive for AI applications since they treat each subset of players as equally important, directly measuring the impact of each player across all possible combinations (see e.g., Equation 2). Additionally, Banzhaf values can be computed more efficiently and tend to be more numerically robust (Karczmarz et al., 2022; Wang & Jia, 2023).
@@
For general set functions, the exact computation of Shapley and Banzhaf values is an NP-hard problem (Deng & Papadimitriou, 1994), so they are often approximated in practice. The problem of estimating Shapley values, especially for model explanation, has been well-studied. A leading method is KernelSHAP (Lundberg & Lee, 2017), a model-agnostic technique that leverages a connection to linear regression, approximating Shapley values by solving a subsampled weighted least squares problem (Charnes et al., 1988; Lundberg & Lee, 2017). KernelSHAP has been further improved with paired sampling (Covert & Lee, 2020) and leverage score sampling Musco & Witter (2024).
@@
In contrast to Shapley values, only a few algorithms have been proposed to compute Banzhaf values for arbitrary set functions. For tree-structured set functions, such as decision tree based models, exact Banzhaf values can be efficiently computed (Karczmarz et al., 2022). For general set functions (e.g., opaque neural networks), Monte Carlo sampling can be used to estimate each Banzhaf value separately (Bachrach et al., 2010). The Maximum Sample Reuse (MSR) algorithm reuses samples for the estimates of different Banzhaf values (Wang & Jia, 2023). However, MSR estimates are correlated and have larger magnitude, and hence high variance as discussed in Section 4.1.
In this work, we apply ideas that have been proven effective for estimating Shapley values to Banzhaf values (Lundberg & Lee, 2017; Covert & Lee, 2020; Musco & Witter, 2024). Our starting point is a formulation of Banzhaf values in terms of a specific linear regression problem (Section 3.1). While a connection to linear regression was known for ‘simple’ set functions (i.e., set functions with binary outputs that satisfy a monotonicity property) in Hammer & Holzman (1992), to the best of our knowledge, we are the first to establish the connection for general set functions. We leverage this connection to design a new algorithm for approximating Banzhaf values: we propose Kernel Banzhaf, inspired by KernelSHAP, to solve a subsampled instance of the Banzhaf linear regression problem (Section 3.2). Similar to Leverage SHAP (Musco & Witter, 2024), we exploit leverage score sampling to give approximation guarantees on the $\ell_{2}$-norm error.
We show that with $O(n\log\frac{n}{\delta}+\frac{n}{\delta\epsilon})$ samples, Kernel Banzhaf produces estimated Banzhaf values $\hat{\bm{\phi}}$ such that $\|\hat{\bm{\phi}}-\bm{\phi}\|_{2}^{2}\leq\epsilon\gamma\|\bm{\phi}\|_{2}^{2}$ with probability $1-\delta$, where $\bm{\phi}$ are the exact Banzhaf values.111 We believe that the $\gamma$ parameter is necessary for linear regression-based algorithms (see Section 3.3).  In contrast, Wang & Jia (2023) show that with $O(\frac{n}{\epsilon}\log\frac{n}{\delta})$ samples, MSR produces estimates such that $\|\hat{\bm{\phi}}^{\textrm{MSR}}-\bm{\phi}\|_{2}^{2}\leq\epsilon$. However, their result assumes that the set function output is restricted to $[0,1]$, excluding regression and generative AI tasks.222 Because the upper bound of their guarantee has no dependence on $\|\bm{\phi}\|_{2}^{2}$, their assumption on the range of the set function effectively normalizes the set function to satisfy the guarantee.
@@
Figure 1: Comparison of exact and estimated Banzhaf values across datasets and estimators using $20n$ samples. Each subplot, labeled with its estimator, shows the normalized estimated versus exact Banzhaf values across all features for a randomly selected explicand from each dataset. Points closer to the diagonal line indicate more accurate estimates; the qualitative evaluation and the $\ell_{2}$-norm error indicate that Kernel Banzhaf is more accurate than the MC and MSR estimators.
Beyond theoretical guarantees, we extensively evaluate the empirical performance of Kernel Banzhaf. Our experimental results show that Kernel Banzhaf produces estimates which are closer to the true Banzhaf values than MC and MSR (Figure 1). In prior work, estimators for Banzhaf values on large datasets were compared in terms of their convergence (Wang & Jia, 2023). Although convergence is an important property, it does not guarantee accuracy, particularly if the estimators have not been proven to be unbiased. In our work, we evaluate the accuracy of Banzhaf estimators for large datasets by using the algorithms of Karczmarz et al. (2022) to compare against exact values. Our findings, across eight datasets, suggest that Kernel Banzhaf is more robust and accurate than prior estimators. Additionally, we compare Kernel Banzhaf with state-of-the-art KernelSHAP algorithms and empirically demonstrate that in semivalue estimation, Kernel Banzhaf provides robustness across all scales and greater efficiency particularly when dealing with a large number of features. We further substantiate this finding by examining the condition number of these regression-based methods.
Contributions
@@
    We propose Kernel Banzhaf, a regression-based approximation algorithm for estimating Banzhaf values for general set functions.
@@
    We show that Kernel Banzhaf provably returns accurate solutions. Our analysis requires non-trivial modifications to the standard leverage score sampling analysis. We argue that, up to log factors and the dependence on $\epsilon$, our analysis is the best possible.
@@
    We run a detailed experimental evaluation using eight popular datasets to evaluate the performance of algorithms for approximating Banzhaf values and Shapley values in feature attribution tasks, varying the number of samples and the reliability of the set function to simulate practical conditions. Unlike prior work, our experiments evaluate estimators relative to the true Banzhaf values, rather than relying just on convergence metrics. Our findings demonstrate that Kernel Banzhaf significantly outperforms existing methods in efficiency and robustness.
@@ 2 Background
Let $n$ be the number of players and $v:2^{[n]}\to\mathbb{R}$ be a set function. The Shapley value of player $i$ for $i\in[n]\vcentcolon=\{1,\ldots,n\}$ is
$\phi^{\textnormal{Shapley}}_{i}=\frac{1}{n}\sum_{S\subseteq[n]\setminus\{i\}}\binom{n-1}{|S|}^{-1}[v(S\cup\{i\})-v(S)]$ (1)
while the Banzhaf value is
$\phi^{\textnormal{Banzhaf}}_{i}=\frac{1}{2^{n-1}}\sum_{S\subseteq[n]\setminus\{i\}}[v(S\cup\{i\})-v(S)]$ (2)
Shapley values and Banzhaf values each uniquely satisfy four different properties. We defer the description of these properties to Appendix H. By default, we shall use $\phi$ to denote $\phi^{\textnormal{Banzhaf}}$ unless explicitly stated otherwise and use $\bm{\phi}\in\mathbb{R}^{n}$ to denote the vector $[\phi_{1},\ldots,\phi_{n}]$.
We can recover various tasks in machine learning by choosing different ways of defining the set function. In this work, we will consider one of the most popular tasks in explainable AI: feature attribution (also known as feature importance and feature influence). Let $M$ be some model which takes in input $\mathbf{x}\in\mathbb{R}^{n}$. In the version we consider, the feature attribution seeks to explain the predictions of the model on an explicand observation $\mathbf{x}^{e}$, which is the target data point for which we want to understand the contributions of individual features. Given a subset of features $S$, define $\mathbf{x}^{S}$ as the observation where $\mathbf{x}^{S}_{i}=\mathbf{x}^{e}_{i}$ if feature $i\in S$ and, otherwise, $\mathbf{x}^{\bar{S}}$ is sampled from a distribution, as discussed in Appendix B. $v(S)=\mathbb{E}[M(\mathbf{x}^{S})]$, where the expectation is over the sampling of the features $i\notin S$. We can adapt the algorithm to other machine learning tasks such as feature (or data) selection by redefining $v(S)$ as the loss of a model trained using only the features (or observations) within subset $S$.
@@ 3 Kernel Banzhaf
The starting point of our work is a formulation of Banzhaf values in terms of a specially structured linear regression problem, and in Theorem 3.2, we show that Banzhaf values are the exact solution to this linear regression problem. Then we propose the Kernel Banzhaf algorithm for estimating Banzhaf values.
@@ 3.1 Linear Regression Formulation
Consider binary vectors $\mathbf{z}\in\{-1,1\}^{n}$. In a slight abuse of notation, we will use $\mathbf{z}$ as input to the set function where the corresponding set is induced by $\mathbf{z}$. We will also use $\mathbf{z}$ as an index for the following matrices.
  • Let $\mathbf{A}\in\mathbb{R}^{2^{n}\times n}$ where the row corresponding to $\mathbf{z}$ is given by $[\mathbf{A}]_{\mathbf{z}}=\frac{1}{2}\mathbf{z}^{\top}$.
  • Let $\mathbf{b}\in\mathbb{R}^{2^{n}}$ where the entry corresponding to $\mathbf{z}$ is given by $\mathbf{b}_{\mathbf{z}}=v(\mathbf{z})$.
The following observation about $\mathbf{A}$ will be helpful in our analysis.
Observation 3.1.
$\mathbf{A}^{\top}\mathbf{A}=2^{n-2}\mathbf{I}$
Proof.
The $(i,j)$ entry in $\mathbf{A^{\top}A}$ is given by
$[\mathbf{A^{\top}A}]_{i,j}=\sum_{\mathbf{z}}\frac{1}{4}z_{i}z_{j}$ (3)
If $i\neq j$, there are $2^{n-1}$ terms of $-\frac{1}{4}$ when $z_{i}\neq z_{j}$ and $2^{n-1}$ terms of $+\frac{1}{4}$ when $z_{i}=z_{j}$ hence Equation 3 is 0. If $i=j$, we have $2^{n}$ terms of $+\frac{1}{4}$ hence Equation 3 is $2^{n-2}$. Together, this gives that $\mathbf{A^{\top}A}=2^{n-2}\mathbf{I}$. ∎
With this observation, we will establish that Banzhaf values are the solution to the linear regression problem defined on $\mathbf{A}$ and $\mathbf{b}$. A similar result was known but only for the case where $v$ is a simple set function (the output is binary $\{0,1\}$ and the function satisfies a monotonicity property) (Hammer & Holzman, 1992).
@@
The Banzhaf values are the solution to the linear regression problem induced by $\mathbf{A}$ and $\mathbf{b}$, i.e.,
$\bm{\phi}=\operatorname*{arg\,min}_{\mathbf{x}}\|\mathbf{Ax-b}\|_{2}.$ (4)
The proof shows that the optimal linear regression solution is equal to the Banzhaf values.
Proof.
By setting the gradient of the objective function to $\mathbf{0}$, we have that
$\operatorname*{arg\,min}_{\mathbf{x}}\|\mathbf{Ax-b}\|_{2}=(\mathbf{A^{\top}A})^{-1}\mathbf{A}^{\top}\mathbf{b}.$ (5)
We will analyze the right hand side. By Observation 3.1, we have $\mathbf{A^{\top}A}=2^{n-2}\mathbf{I}$. Then $(\mathbf{A^{\top}A})^{-1}=\frac{1}{2^{n-2}}\mathbf{I}$. Continuing, we have
$(\mathbf{A^{\top}A})^{-1}\mathbf{A^{\top}b}=\frac{1}{2^{n-2}}\mathbf{A^{\top}b}=\frac{1}{2^{n-2}}\sum_{\mathbf{z}}\frac{1}{2}\mathbf{z}v(\mathbf{z})=\frac{1}{2^{n-1}}\sum_{\mathbf{z}}\mathbf{z}v(\mathbf{z}).$ (6)
Now consider entry $i$ given by
$[(\mathbf{A^{\top}A})^{-1}\mathbf{A^{\top}b}]_{i}=\frac{1}{2^{n-1}}\sum_{\mathbf{z}}z_{i}v(\mathbf{z})=\frac{1}{2^{n-1}}\sum_{S\subseteq[n]\setminus\{i\}}v(S\cup\{i\})-v(S)$ (7)

### v2

@@ Abstract
Banzhaf values provide a popular, interpretable alternative to the widely-used Shapley values for quantifying the importance of features in machine learning models. Like Shapley values, computing Banzhaf values exactly requires time exponential in the number of features, necessitating the use of efficient estimators. Existing estimators, however, are limited to Monte Carlo sampling methods. In this work, we introduce Kernel Banzhaf, the first regression-based estimator for Banzhaf values. Our approach leverages a novel regression formulation, whose exact solution corresponds to the exact Banzhaf values. Inspired by the success of Kernel SHAP for Shapley values, Kernel Banzhaf efficiently solves a sampled instance of this regression problem. Through empirical evaluations across eight datasets, we find that Kernel Banzhaf significantly outperforms existing Monte Carlo methods in terms of accuracy, sample efficiency, robustness to noise, and feature ranking recovery. Finally, we complement our experimental evaluation with strong theoretical guarantees on Kernel Banzhaf’s performance.
Machine Learning, ICML
@@ 1 Introduction
The increasing complexity of AI models has intensified challenges associated with model interpretability. Modern machine learning models, such as deep neural networks and ensemble methods, often operate as “opaque boxes.” This opacity makes it difficult for users to understand and trust model predictions, especially in decision-making scenarios like healthcare, finance, and legal applications, which require rigorous justifications. Thus, there is a pressing need for reliable explainability tools to bridge the gap between complex model behaviors and human understanding.
Among the various methods employed within explainable AI, game-theoretic approaches have gained prominence for quantifying the contribution of features both to the overall performance of a machine learning model, and to individual predictions made by the model. The most well-known game-theoretic approach is based on Shapley values, which provide a principled way to attribute the contribution of $n$ individual players to the outcome of a cooperative game, which is defined by a set function $v$ that maps every subset of $\{1,\ldots,n\}$ to a real value (Shapley, 1953).
@@
In the context of feature attribution, each “player” is a feature and $v$ maps a subset of features to, e.g., the overall test loss when the chosen features are used, or the prediction made for a specific individual given access to the chosen features. The Shapley value quantifies the average marginal contribution of a feature on the set function, computed as the weighted average over all possible combinations of features included in the model (Lundberg & Lee, 2017). More formally, the Shapley value $\phi_{i}$ of a player $i\in\{1,\ldots,n\}$ is
$\phi_{i}^{\textnormal{shap}}=\frac{1}{n}\sum_{S\subseteq[n]\setminus\{i\}}\binom{n-1}{|S|}^{-1}[v(S\cup\{i\})-v(S)].$ (1)
While primarily associated with feature attribution (Lundberg & Lee, 2017; Karczmarz et al., 2022), game-theoretic methods like Shapley values have also contributed to broader machine learning tasks beyond explainability, such as feature selection (Covert et al., 2020) and data valuation (Ghorbani & Zou, 2019; Wang & Jia, 2023).
@@
An alternative to Shapley values are Banzhaf values, which also compute each individual’s contribution to an overall outcome (Banzhaf, 1965; Lehrer, 1988). While Shapley values are more widely used, Banzhaf values are often considered more intuitive for AI applications since they treat each subset of players as equally important, directly measuring the impact of each player across all possible combinations. In particular, the Banzhaf value is
$\phi_{i}^{\textnormal{banz}}=\frac{1}{2^{n-1}}\sum_{S\subseteq[n]\setminus\{i\}}[v(S\cup\{i\})-v(S)].$ (2)
In contrast, the Shapley values use non-uniform weight $\binom{n-1}{|S|}^{-1}$, which depends on the subset size. It has also been observed that Banzhaf values tend to be more robust than Shapley values in the context of explainable AI (Karczmarz et al., 2022; Wang & Jia, 2023). These benefits have led to the broad use of Banzhaf values for feature attribution (Datta et al., 2015; Kulynych & Troncoso, 2017; Sliwinski et al., 2018; Patel et al., 2021) and, more recently, for data valuation (Wang & Jia, 2023; Li & Yu, 2024b).
@@
For general set functions, the exact computation of both Shapley and Banzhaf values requires at least $2^{n}$ time, where $n$ is the number of features in our problem. In particular, we must evaluate $v(S)$ for every possible subset $S$ of $\{1,\ldots,n\}$. Accordingly, the problem of approximating Shapley values at lower cost has been widely studied, especially for model explanation. A leading method for Shapley values is Kernel SHAP (Lundberg & Lee, 2017), a model-agnostic technique that leverages a connection to linear regression, approximating Shapley values by solving a sampled weighted least squares problem (Charnes et al., 1988; Lundberg & Lee, 2017). Kernel SHAP has been further improved with paired sampling (Covert & Lee, 2020) and leverage score sampling (Musco & Witter, 2024).
@@
In contrast to Shapley values, only a few algorithms have been proposed to approximate Banzhaf values for arbitrary set functions. For tree-structured set functions, such as those corresponding to decision tree based models, exact Banzhaf values can be efficiently computed (Karczmarz et al., 2022). For general set functions, like those corresponding to neural networks models, Monte Carlo sampling (i.e., estimating the sum in (2) via a subsample) can be used to estimate each Banzhaf value separately (Merrill III, 1982; Bachrach et al., 2010). The Maximum Sample Reuse (MSR) algorithm is a variant of Monte Carlo sampling that reuses samples for the estimates of different Banzhaf values (Wang & Jia, 2023). However, even with reuse, the naive Monte Carlo estimator typically requires a large number of samples (which equates to a large number of evaluations of $v$) to obtaining meaningfully accurate estimates for the Banzhaf values.
Our Contributions.
In this work, we show for the first time that, like Shapley values, Banzhaf values can be approximated far more accurately through approximate regression. To do so, we introduce a novel regression formulation (concretely, a linear least squares regression problem involving $n$ variables) whose solution exactly corresponds to the Banzhaf values (Section 3.1). We then show that this problem can be efficiently solved by subsampling rows, each of which requires just a single evaluation of the set function $v$.
The ultimate payoff is a novel estimator, Kernel Banzhaf, that on all datasets tested, far outperforms existing Monte Carlos estimation methods (see Section 4 for more details). Concretely, for a fixed number of evaluations of $v$, our method typically improves on the accuracy of Banzhaf values estimation by an order of magnitude or more.
While two alternative regression formulations were previously known for Banzhaf values, neither lead to efficient algorithms. One only recovers the values up to an unknown additive shift (Ruiz et al., 1998; Li & Yu, 2024a) and unfortunately, it is unclear how to estimate this shift efficiently. The other exactly recovers the Banzhaf values, but only for binary and monotone set functions (Hammer & Holzman, 1992). Our formulation, in contrast, exactly recovers the Banzhaf values without any assumptions on the set function.
We conduct an extensive experimental evaluation of Kernel Banzhaf in comparison to existing Banzhaf estimators, assessing performance in terms of sample efficiency and robustness to noise. We compare accuracy using both the relative squared $\ell_{2}$-norm error and an objective naturally suggested by our novel regression problem. Furthermore, we evaluate Kernel Banzhaf in a feature ranking task, where the goal is to accurately recover the most important features. Our experiments span diverse model classes, including tree-based models and neural networks. Our findings suggest that Kernel Banzhaf is substantially more accurate and robust than prior estimators (see e.g., Figure 1).
We complement our experimental evaluation of Kernel Banzhaf with strong theoretical guarantees. We prove that, with just $O(n\log n)$ evaluations of $v$, Kernel Banzhaf returns a solution with bounded squared $\ell_{2}$-norm difference to the exact Banzhaf values (see Corollary 3.3). Obtaining similar theoretical guarantees for Monte Carlo sampling and improvements like Maximum Sample Reuse requires making a strong assumption on the maximum magnitude of $v(S)$. See Section 3.3 for further comparison of theoretical results.
@@
Figure 1: Comparison of exact and estimated Banzhaf feature attribution values across eight datasets, where every estimator uses $m=20n$ evaluations of the set function $v$. The evaluations dominate the computational cost of each algorithm. Each subplot, labeled with its estimator, shows normalized estimated versus exact Banzhaf values across all features for a randomly selected data point from each dataset. Points closer to the diagonal line indicate more accurate estimates; the plots suggest that Kernel Banzhaf is more accurate than the Monte Carlo and Maximum Sample Reuse estimators.
@@
    We propose Kernel Banzhaf, the first regression-based approximation algorithm for estimating Banzhaf values for general set functions. The foundation of the algorithm is a novel regression formulation, whose exact solution is the Banzhaf values.
@@
    We show that Kernel Banzhaf provably returns accurate solutions. Our theoretical guarantees depend on the magnitude of the true Banzhaf values rather than the magnitude of the set function, as for prior estimators.
@@
    We extensively evaluate Banzhaf estimator performance across eight popular datasets, in various settings, and under several measures of performance. Our findings demonstrate that Kernel Banzhaf significantly outperforms prior work.
@@ 2 Background
Let $n$ be a positive integer and define $[n]\vcentcolon=\{1,2,\ldots,n\}$. We let $2^{[n]}$ denote the set of all $2^{n}$ possible subsets of $[n]$ (including the emptyset). For set $S$, define the indicator vector $\mathbbm{1}[S]\in\{0,1\}^{n}$ so that $\mathbbm{1}[S]_{i}=1$ if $i\in S$ and 00 otherwise. Let $v:2^{[n]}\to\mathbb{R}$ be a set function. Shapley and Banzhaf values are as defined in Equations 1 and 2 in the prior section. For notational simplicity, we will use $\phi_{i}=\phi_{i}^{\textnormal{banz}}$ in the rest of the paper. Let $\boldsymbol{\phi}\in\mathbb{R}^{n}$ denote the vector $[\phi_{1},\ldots,\phi_{n}]$.
Since we make no structural assumptions about $v$, we consider estimators for computing Banzhaf values that access the function via samples; formally, a single sample corresponds to evaluating $v(S)$ for any $S$ the algorithm chooses.
Banzhaf Values for Feature Attribution. As discussed in the previous section, Shapley and Banzhaf values have a number of applications in machine learning, which correspond to different choices of the set function $v$. For example, we can consider feature selection by defining $v(S)$ as the loss of a model trained only on the features in $S$. Similarly, we can consider data selection by defining $v(S)$ as the loss of a model trained only on the observations in $S$.
One of the most popular applications of Banzhaf values in explainable AI is feature attribution (also known as feature importance and feature influence). Let $\mathcal{M}:\mathbb{R}^{n}\to\mathbb{R}$ be a trained model which takes input $\mathbf{x}\in\mathbb{R}^{n}$. Our goal is to explain the predictions of the model on a particular data point $\mathbf{x}$. Given a subset of features $S$, define $\mathbf{x}^{S}$ as the observation where $\mathbf{x}^{S}_{i}=\mathbf{x}_{i}$ if feature $i\in S$ and, otherwise, $\mathbf{x}^{S}_{i}$ is sampled from the data distribution (possibly conditioned on the features in $S$). Then $v(S)=\mathbb{E}[\mathcal{M}(\mathbf{x}^{S})]$, where the expectation is over the sampling of the features $i\notin S$. This expectation can be estimated via sampling, although for tree-based models, it is actually possible to exactly compute the Banzhaf feature attribution values in an efficient way (Karczmarz et al., 2022). This is valuable in evaluating the performance of Banzhaf value approximation methods since, for these models, we can obtain the true Banzhaf values to compare against.
@@ 3 Kernel Banzhaf
The starting point of our work is a formulation of Banzhaf values in terms of a structured linear regression problem with an exponential number of rows. We show in Theorem 3.1 that Banzhaf values are the exact solution to this linear regression problem. Our efficient Kernel Banzhaf algorithm is then obtained by solving the regression problem approximately using just a small subset of rows.
@@ 3.1 Linear Regression Formulation
Let $\mathbf{A}\in\mathbb{R}^{2^{n}\times n}$ be a design matrix and $\mathbf{b}\in\mathbb{R}^{2^{n}}$ be a target vector. We will use subsets $S\subseteq[n]$ to index the rows of $\mathbf{A}$ and entries of $\mathbf{b}$. In particular, $[\mathbf{A}]_{S,i}$ denotes the $i^{\text{th}}$ entry of row $S$. We set:
$[\mathbf{A}]_{S,i}=\begin{cases}+\frac{1}{2}&\textnormal{ if $i\in S$}\\ -\frac{1}{2}&\textnormal{ if $i\notin S$}\end{cases}$ (3)
and
$\mathbf{b}_{S}=v(S).$ (4)
@@
Consider $\mathbf{A}$ and $\mathbf{b}$ as defined in Equations 3 and 4, respectively. Let
$\mathbf{x}^{*}=\operatorname*{arg\,min}_{\mathbf{x}}\|\mathbf{Ax-b}\|_{2}^{2}.$
Then $\boldsymbol{\phi}=\mathbf{x}^{*}$, where $\boldsymbol{\phi}$ are the Banzhaf values of $v$.
Proof of Theorem 3.1.
As is standard, we have that:
$\operatorname*{arg\,min}_{\mathbf{x}}\|\mathbf{Ax-b}\|_{2}^{2}=(\mathbf{A^{\top}A})^{-1}\mathbf{A}^{\top}\mathbf{b}.$
We will analyze the right hand side. The $(i,j)$ entry in $\mathbf{A^{\top}A}$ is given by
$[\mathbf{A^{\top}A}]_{i,j}=\sum_{S\subseteq[n]}\left(\mathbbm{1}[i\in S]-\frac{1}{2}\right)\left(\mathbbm{1}[j\in S]-\frac{1}{2}\right)$ (5)
If $i\neq j$, there are $2^{n-1}$ terms of $-\frac{1}{4}$ when $\mathbbm{1}[i\in S]\neq\mathbbm{1}[j\in S]$ and $2^{n-1}$ terms of $+\frac{1}{4}$ when $\mathbbm{1}[i\in S]=\mathbbm{1}[j\in S]$, hence Equation 5 is 0. If $i=j$, we have $2^{n}$ terms of $+\frac{1}{4}$ hence Equation 5 is $2^{n-2}$. Together, this gives that
$\mathbf{A^{\top}A}=2^{n-2}\mathbf{I}.$ (6)
Then $(\mathbf{A^{\top}A})^{-1}=\frac{1}{2^{n-2}}\mathbf{I}$. Continuing, we have
$(\mathbf{A^{\top}A})^{-1}\mathbf{A^{\top}b}$ $=\frac{1}{2^{n-2}}\mathbf{A^{\top}b}$
$=\frac{1}{2^{n-2}}\sum_{S\subseteq[n]}\left(\mathbbm{1}[S]-\frac{1}{2}\mathbf{1}\right)v(S).$
We can write the $i^{\text{th}}$ entry of the above as:
$[(\mathbf{A^{\top}A})^{-1}\mathbf{A^{\top}b}]_{i}$ $=\frac{1}{2^{n-2}}\sum_{S\subseteq[n]}\left(\mathbbm{1}[i\in S]-\frac{1}{2}\right)v(S)$
$=\frac{1}{2^{n-1}}\sum_{S\subseteq[n]\setminus\{i\}}v(S\cup\{i\})-v(S)$
@@
Other Regression Formulations.
Despite the simplicity of the regression problem above, we are not aware of any prior work that considers it. As discussed, a regression formulation for Banzhaf values is known for the special case where $v$ is binary and monotone i.e., $v:2^{[n]}\to\{0,1\}$ and $v(S)\leq v(T)$ for $S\subseteq T$ (Hammer & Holzman, 1992). Another relevant regression problem has the property that $\boldsymbol{\phi}=\mathbf{x}^{*}+c\mathbf{1}$ for $\mathbf{1}\in\mathbb{R}^{n}$ the all-ones vector and $c$ an unknown number that depends on $v$ (Ruiz et al., 1998; Li & Yu, 2024a); however, it is not clear how to estimate $c$ efficiently and thus recover the Banzhaf values.